OpenStudy (anonymous):

RATIONALIZE THIS:

OpenStudy (anonymous):

\[\frac{ \sqrt[3]{x + 24} - 3 }{ \sqrt{x + 1} - 2 }\]

terenzreignz (terenzreignz):

I assume what you want is to get rid of the radical on the denominator?

OpenStudy (anonymous):

so sorry for the late reply. but yes.

terenzreignz (terenzreignz):

Well, the trick I use here basically involves 'not caring about the numerator' Put it this way: Generic fraction with a radical on the denominator... \[\Large \frac{N}{\sqrt{r}\pm k}\]

terenzreignz (terenzreignz):

N is the numerator (which I don't care about at the moment) and we have \(\large \sqrt{r}\pm k\) in the denominator... and we want to remove the radical... catch me so far?

OpenStudy (anonymous):

yep. i tried that. actually I've reached this\[\frac{ \sqrt[3]{x+24-3} }{\sqrt{x+1}-2} \times \frac{ \sqrt[3]{x+24+3} }{\sqrt{x+1}+2} \times \frac{ \sqrt{x+1}+2 }{\sqrt{x+1}+2 }\]

OpenStudy (anonymous):

yep. i understand you :)

terenzreignz (terenzreignz):

Oh wait... something's wrong...

terenzreignz (terenzreignz):

YOU paid attention to the numerator... I told you not to :P All you do is multiply both the numerator and denominator by the *conjugate* of the denominator... (fancy word, but all it means is that you change the - to a +)

OpenStudy (anonymous):

what? omg. sorry. we were taught to do that in school. Haha

OpenStudy (anonymous):

its all about finding the limits. my answer was indeterminate so now, I had to like rationalize it or something

terenzreignz (terenzreignz):

Oh... okay... then honestly, I don't know how to go about this... (something specific to your school instruction?)

terenzreignz (terenzreignz):

Could you post the original question? :3

OpenStudy (radar):